OpenStudy (anonymous):

solve the formula for the specified variable: A=1/2bh for h

OpenStudy (anonymous):

Is it \[A={1 \over 2bh}?\]

OpenStudy (anonymous):

your solving for h

OpenStudy (anonymous):

I am asking if what I wrote above is the original expression?

OpenStudy (anonymous):

If so, then: \[(2Ab)h=1 \implies h={1 \over 2Ab}\]

OpenStudy (anonymous):

its A=1/2(one half)bh

OpenStudy (anonymous):

But if the original expression is: \[A={1 \over 2}bh\] then: \[h={2A \over b}\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

np.

OpenStudy (anonymous):

can you try and help me with some other?!

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

the total cost in dollars for a certain company to produce x empty jars to be used by a jelly producer is given by the function C(x)=.07x=12,000. find C(90,000), the cost of producing 90,000 jars

OpenStudy (anonymous):

Is it C(x)=0.7x-12,000?

OpenStudy (anonymous):

A.) $63,000 B.) $75,000 C.) $12.70 D.) $90,012

OpenStudy (anonymous):

I don't want choices. You made a mistake in writing the formula. I want you to check it. Is +12,000 or -12,000?

OpenStudy (anonymous):

positive 12,000

OpenStudy (anonymous):

If so, then the cost of 90,000 jars is: \[C(90,000)=0.7(90,000)+12,000=$75,000\]

OpenStudy (anonymous):

sorry for the misunderstanding thanks so much!

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

You could give me a medal :)